Suppose Set B contains 61 elements and the total number elements in either Set A or Set B is 112. If the Sets A and B have 38 elements in common, how many elements are contained in set A?
Found 3 solutions by stanbon, ikleyn, Theo:
Answer by stanbon(75887)

Suppose Set B contains 61 elements and the total number elements in either Set A or Set B is 112. If the Sets A and B have 38 elements in common, how many elements are contained in set A?
-----
A + B - (A and B) = (A or B)
-----
A + 61 - 38 = 112
------
A = 150-61
----
A = 89 elements
---------------------
Cheers,
Stan H.
----------------

Answer by ikleyn(52874)
Answer by Theo(13342)

set B contains 61 elements.
total number of elements in set A or set B is 112.
sets A and B have 38 element in common.

the formula is number of elements in set (A or B) is equal to the numbers of elements in set A plus the number of elements in set B minus the elements that are in both set A and set B.

the formula assumes that the elements in set A include the elements that are also in set B, and it assumes that the elements in set B include the elements that are also in set A.

because of this, the elements in set A and B are double counted, once as part of set A and again as part of set B.

the subtraction of the number of elements in both set A and B eliminates the double counting.

you get:

112 = 61 + A - 38

A represents the number of elements in set A

solve for A to get A = 112 - 61 + 38 = 89

there are 61 elements in set A and 89 elements in set B.

each set is assumed to contain elements that are in both.

take out the elements in set A that are also in set B.
take out the elements in set B that are also in set A.

you get the number of elements in set A only are 61 - 38 = 23.
you get the number of elements in set B only are 89 - 38 = 51.

now count the number of elements that are in both set A and B separately.

you get 23 in A and 51 in B and 38 in both for a total of 112.